# Distributional Diffusion Models with Scoring Rules

### v2

@@ 6.3 Robotics experiments
Experimental setup. We demonstrate here the effectiveness of our distributional diffusion models for robotics applications. We experiment with diffusion policies on the Libero (Liu et al., 2024) benchmark, a life-long learning benchmark that consists of 130 language-conditioned robotic manipulation tasks. In our experiments we focus on Libero-Long, the most challenging suite with 10 tasks that features long-horizon tasks with diverse object interactions. Our multi-task diffusion policy is conditioned on the encoded visual and proprioceptive observations, and the task descriptions and generates a chunk of 8 7-dimensional actions to execute; see Section I.3 for additional details on the architecture.
@@
Results. Similar to the results on image experiments, $\lambda=0$ works best when more diffusion steps $(50)$ are used during evaluation. Using $\lambda>0$ gives best results when fewer diffusion steps are used $(2,16)$. We note that the performance slightly deteriorates with $\lambda>0$ when the number of diffusion steps is large, see Figure 9.
@@ Appendix A Notation
Name Notation Definition First appearance
$\cdot$ $\rho$ Continuous negative definite kernel Section 2.2
$\cdot$ $k$ Continuous positive definite kernel Section 2.2
Scoring rule $S(p,y)$cdot Section 2.2
Expected score $S(p,q)$ $S(p,q)=\mathbb{E}_{q}[S(p,Y)]$ Section 2.2
Kernel score $S_{\rho}(p,y)$ $S_{\rho}(p,y)=\frac{1}{2}\mathbb{E}[\rho(X,X^{\prime})]-\mathbb{E}[\rho(X,y)]$ Section 2.2
Energy score $S_{\beta}(p,y)$ $S_{\beta}(p,y)=S_{\rho}(p,y)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Section 2.2
Expected Kernel score $S_{\rho}(p,q)$ $S_{\rho}(p,q)=\mathbb{E}_{q}[S_{\rho}(p,Y)]$ Section 2.2
Expected Energy score $S_{\beta}(p,q)$ $S_{\beta}(p,q)=S_{\rho}(p,q)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Section 2.2
Generalized Kernel score $S_{\lambda,\rho}(p,y)$ $S_{\lambda,\rho}(p,y)=\frac{\lambda}{2}\mathbb{E}_{p\otimes p}[\rho(X,X^{\prime})]-\mathbb{E}[\rho(X,y)]$ Section 2.2
Generalized Energy score $S_{\lambda,\beta}(p,y)$ $S_{\lambda,\beta}(p,y)=S_{\lambda,\rho}(p,y)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Section 2.2
Expected Generalized Kernel score $S_{\lambda,\rho}(p,q)$ $S_{\lambda,\rho}(p,q)=\mathbb{E}_{q}[S_{\lambda,\rho}(p,Y)]$ Section 2.2
Expected Generalized Energy score $S_{\lambda,\beta}(p,q)$ $S_{\lambda,\beta}(p,q)=S_{\lambda,\rho}(p,q)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Section 2.2
Squared Maximum Mean Discrepancy $D_{\rho}(p,q)$ $D_{\rho}(p,q)=S_{\rho}(q,q)-S_{\rho}(p,q)$ Section 2.2
Energy Distance $D_{\beta}(p,q)$ $D_{\beta}(p,q)=D_{\rho}(p,q)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Section 2.2
Diffusion Loss $\mathcal{L}_{\text{diff}}(\theta)$ $\mathcal{L}_{\text{diff}}(\theta)=\int^{1}_{0}w_{t}\mathbb{E}_{p_{0,t}}\left[||X_{0}-\hat{x}_{\theta}(t,X_{t})||^{2}\right]\mathrm{d}t$ Section 2.1
Energy Diffusion Loss $\mathcal{L}(\theta)$ $\mathcal{L}(\theta)=-\int_{0}^{1}w_{t}\mathbb{E}_{p_{0,t}}\left[S_{\lambda,\beta}(p_{0|t}^{\theta}(\cdot|X_{t}),X_{0})\right]\mathrm{d}t$ Section 3
Kernel Diffusion Loss $\mathcal{L}(\theta)$ $\mathcal{L}(\theta)=-\int_{0}^{1}w_{t}\mathbb{E}_{p_{0,t}}\left[S_{\lambda,\rho}(p_{0|t}^{\theta}(\cdot|X_{t}),X_{0})\right]\mathrm{d}t$ Section 3
Joint Energy Diffusion Loss $\mathcal{L}_{\mathrm{joint}}(\theta)$ $\mathcal{L}_{\mathrm{joint}}(\theta)=-\int_{0}^{1}w_{t}\mathbb{E}_{p_{0,t}}\left[S_{\lambda,\beta}(p_{0|t}^{\theta}p_{t},(X_{0},X_{t}))\right]\mathrm{d}t$ Section 4.2
@@ Appendix B Correspondence between Discrepancy and Diffusion Loss
Proposition B.1: Assume that $D_{\rho}(p,q)=\|\mathbb{E}_{p}[X]-\mathbb{E}_{q}[X]\|^{2}$ and that $p=\updelta_{\hat{x}_{\theta}(t,X_{t})}$ and $q=p_{0|t}(\cdot|X_{t})$. Then, there exists a constant $C\in\mathbb{R}$, independent of $\theta$, such that $\mathbb{E}_{p_{t}}[D_{\rho}(p,q)]=\mathbb{E}_{p_{0,t}}[\|X_{0}-\hat{x}_{\theta}(t,X_{t})\|^{2}]+C,$ (25) with $C=-\mathbb{E}_{p_{0,t}}[\|X_{0}-\mathbb{E}[X_{0}|X_{t}]\|^{2}]$.
@@ D.1 Energy diffusion loss and Maximum Mean Discrepancy diffusion loss
In this section, we first compare the energy diffusion loss (14) and another loss based on Maximum Mean Discrepancy (MMD), denoted MMD diffusion loss. We show that these two losses are equal up to a constant which does not depend on the network parameters $\theta$. However, this additional term appearing in the MMD diffusion loss is untractable, even though it could be potentially estimated using important sampling techniques. We consider the case where the kernel $k$ is given by $k(x,x^{\prime})=-\|x-x^{\prime}\|$. While our discussion can be extended to other settings, we also focus on the case where $\lambda=1,\beta=1$ and $w_{t}=1$ for simplicity.
@@ D.2 Joint Diffusion Energy Loss
We recall that the term which does not depend on $\theta$ in $\mathcal{L}_{\mathrm{MMD}}$ is difficult to approximate, the proposed approximation using importance sampling (74) is indeed expected to perform poorly. On the contrary the term which does not depend on $\theta$ in $\mathcal{L}_{\mathrm{MMD},\mathrm{joint}}$ is very easy to approximate. This suggests that if one would optimize the loss with respect to the parameters of the kernel, then the choice $\mathcal{L}_{\mathrm{MMD},\mathrm{joint}}$ is more suitable. However, in the case where the parameters are fixed, as it is the case in our framework, then we will show in Section D.5 that $\mathcal{L}_{\mathrm{MMD}}$ is more sample efficient.
@@ D.5 Sample efficiency
Appendix E Diffusion compatible kernels
@@
$k_{\texttt{rbf}}(x,x^{\prime})$ $=\mathrm{exp}\left[-\|x-x^{\prime}\|^{2}/\sigma\right],$ (122b)
@@ Appendix F DDIM updates from a Stochastic Differential Equation perspective
We first show that difference choices of forward process yields different interpolation densities $p_{s|0,t}$ in Section F.1. With a careful choice of the forward process and appropriate parameters, we show that we can recover DDIM updates (Song et al., 2021a) in Section F.2.
@@ F.1 General forward processes
Interpolation and forward process.
Consider the following interpolation
$\mathbf{X}_{t}=\alpha_{t}\mathbf{X}_{0}+\sigma_{t}\mathbf{Z},\qquad\mathbf{X}_{0}\sim p_{0}.$ (140)
where $\mathbf{Z}\sim\mathcal{N}(0,\mathrm{Id})$ and the equality holds in distribution. Next, we associate a forward process with this interpolation. First, note that one possible process is given by $\mathbf{X}_{0}\sim p_{0}$ and
$\mathrm{d}\mathbf{X}_{t}$ $=[\partial_{t}\alpha_{t}\mathbf{X}_{0}+\partial_{t}\sigma_{t}\mathbf{Z}]\mathrm{d}t$ (141)
$=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\left\{\partial_{t}\sigma_{t}-\partial_{t}\alpha_{t}\frac{\sigma_{t}}{\alpha_{t}}\right\}\mathbf{Z}\right]\mathrm{d}t,$ (142)
where we have used (140). In addition, one can show that another valid sampler is given by
$\mathrm{d}\mathbf{X}_{t}=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\left\{\partial_{t}\sigma_{t}-\partial_{t}\alpha_{t}\frac{\sigma_{t}}{\alpha_{t}}\right\}((1-\varepsilon^{2})\mathbf{Z}+\varepsilon^{2}\mathbb{E}[\mathbf{Z}|\mathbf{X}_{t}])\right]\mathrm{d}t,$ (143)
with $\varepsilon\in[0,1]$. Note that this is a mean-field equation, in the sense that (143) does not depend only on $\mathbf{X}_{t}$ but also on its distribution. More precisely, we have the following result.
Proposition F.1: Let $\varepsilon\in[0,1]$. Let $\mathbf{Z}\sim\mathcal{N}(0,\mathrm{Id})$ and $\mathbf{X}_{0}\sim p_{0}$ and assume that the following mean-field ODE admits a solution $\mathrm{d}\mathbf{X}_{t}=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\left\{\partial_{t}\sigma_{t}-\partial_{t}\alpha_{t}\frac{\sigma_{t}}{\alpha_{t}}\right\}\left((1-\varepsilon_{t}^{2})z+\varepsilon_{t}^{2}\frac{\int_{\mathbb{R}^{d}}\tilde{z}p_{t}(\mathbf{X}_{t}|\tilde{z})p(\tilde{z})r\mathrm{d}\tilde{z}}{\int_{\mathbb{R}^{d}}p_{t}(\mathbf{X}_{t}|\tilde{z})p(\tilde{z})\mathrm{d}\tilde{z}}\right)\right]\mathrm{d}t,$ (144) where $p(z)=\mathcal{N}(z;0,\mathrm{Id})$ and $p_{t}(\cdot|z)$ is the density of $\mathbf{X}_{t}$ conditionally to $\mathbf{Z}=z$. Denote $p_{t}(x_{t})=\int_{\mathbb{R}^{d}}p_{t}(x_{t}|\tilde{z})p(\tilde{z})\mathrm{d}\tilde{z}$, where $p(\tilde{z})=\mathcal{N}(\tilde{z};0,\mathrm{Id})$. Then, we have that for any $t\in[0,1]$, $\alpha_{t}\mathbf{X}_{0}+\sigma_{t}\mathbf{Z}\sim p_{t}$.
Now, we recall that using Tweedie’s identity we have
$\nabla\log p_{t}(x_{t})=\frac{1}{\sigma_{t}^{2}}\mathbb{E}[\alpha_{t}\mathbf{X}_{0}-\mathbf{X}_{t}|\mathbf{X}_{t}=x_{t}]=-\frac{1}{\sigma_{t}}\mathbb{E}[\mathbf{Z}|\mathbf{X}_{t}=x_{t}].$ (145)
The following non-Markov diffusion admits the same marginals as (144)
$\mathrm{d}\mathbf{X}_{t}$ $=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\left\{\partial_{t}\sigma_{t}-\partial_{t}\alpha_{t}\frac{\sigma_{t}}{\alpha_{t}}\right\}((1-\varepsilon^{2})\mathbf{Z}+\varepsilon^{2}\mathbb{E}[\mathbf{Z}|\mathbf{X}_{t}])+\frac{\varepsilon_{t}^{2}}{2}\nabla\log p_{t}(\mathbf{X}_{t})\right]\mathrm{d}t+\varepsilon_{t}\mathrm{d}\mathbf{B}_{t}$ (146)
$=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\left\{\partial_{t}\sigma_{t}-\partial_{t}\alpha_{t}\frac{\sigma_{t}}{\alpha_{t}}\right\}((1-\varepsilon_{t}^{2})\mathbf{Z}+\varepsilon^{2}\mathbb{E}[\mathbf{Z}|\mathbf{X}_{t}])-\frac{\varepsilon_{t}^{2}}{2\sigma_{t}}\mathbb{E}[\mathbf{Z}|\mathbf{X}_{t}]\right]\mathrm{d}t+\varepsilon_{t}\mathrm{d}\mathbf{B}_{t},$ (147)
as we have simply added a Langevin term of invariant distribution $p_{t}$ to (143). Next, we define
@@
In that case, (146) simplifies to
$\mathrm{d}\mathbf{X}_{t}=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\frac{g_{t}^{2}}{2\sigma_{t}}(1-\varepsilon^{2})\mathbf{Z}\right]\mathrm{d}t+\varepsilon g_{t}\mathrm{d}\mathbf{B}_{t}.$ (149)
In (149), we identify different forward processes which are all non-Markov except in the case where $\varepsilon=1$. The remarkable property of these forward processes is that they all recover the interpolation (140). This means that the forward trajectories might be very different but they all admit the same marginal distributions. More formally, we get the following result.
Proposition F.2: For any $t\in[0,1]$, let $\varepsilon_{t}\in[0,1]$. Let $\mathbf{Z}\sim\mathcal{N}(0,\mathrm{Id})$ and $\mathbf{X}_{0}\sim p_{0}$. Additionally, assume that the following SDE admits a solution $\mathrm{d}\mathbf{X}_{t}=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\frac{g_{t}^{2}}{2\sigma_{t}}(1-\varepsilon^{2})\mathbf{Z}\right]\mathrm{d}t+\varepsilon g_{t}\mathrm{d}\mathbf{B}_{t},$ (150) where $(\mathbf{B}_{t})_{t\in[0,1]}$ is a $d$-dimensional Brownian motion, $p(z)=\mathcal{N}(z;0,\mathrm{Id})$ and $p_{t}(\cdot|z)$ is the density of $\mathbf{X}_{t}$ conditionally to $\mathbf{Z}=z$. Denote $p_{t}(x_{t})=\int_{\mathbb{R}^{d}}p_{t}(x_{t}|z)p(z)\mathrm{d}z$. We have that for any $t\in[0,1]$, $\alpha_{t}\mathbf{X}_{0}+\sigma_{t}\mathbf{Z}\sim p_{t}$.
@@
$\mathbf{X}_{t}=\frac{\alpha_{t}}{\alpha_{s}}\mathbf{X}_{s}+(1-\varepsilon^{2})\left(\sigma_{t}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}\right)\mathbf{X}_{1}+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}},$ (151)
@@
$\mathbf{X}_{t}-\alpha_{t}\mathbf{X}_{0}$ $=-\alpha_{t}\mathbf{X}_{0}+\frac{\alpha_{t}}{\alpha_{s}}\mathbf{X}_{s}+(1-\varepsilon^{2})\left(\sigma_{t}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}\right)\frac{\mathbf{X}_{s}-\alpha_{s}\mathbf{X}_{0}}{\sigma_{s}}+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}}$ (152)
$=\left[\varepsilon^{2}\frac{\alpha_{t}}{\alpha_{s}}+(1-\varepsilon^{2})\frac{\sigma_{t}}{\sigma_{s}}\right](\mathbf{X}_{s}-\alpha_{s}\mathbf{X}_{0})+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}}.$ (153)
We can now compute the probability distribution $p_{s,t|0}$. This will allow us to compute $p_{s|0,t}$ using Gaussian posteriors. Using (153), we have

### v3

@@ 6.3 Robotics experiments
Experimental setup. We demonstrate here the effectiveness of our distributional diffusion models for robotics applications. We experiment with diffusion policies on the Libero (Liu et al., 2024) benchmark, a life-long learning benchmark that consists of 130 language-conditioned robotic manipulation tasks. In our experiments, we used 4 Libero suites with 10 tasks in each: Libero-Long, Libero-Goal, Libero-Object, and Libero-Spatial. Our multi-task diffusion policy is conditioned on the encoded visual and proprioceptive observations, and the task descriptions and generates a chunk of 8 7-dimensional actions to execute; see Section I.3 for additional details on the architecture.
@@
Results. For the main result, we focus on Libero-Long, the most challenging suite with 10 tasks that features long-horizon tasks with diverse object interactions. Similar to the results on image experiments, $\lambda=0$ works best when more diffusion steps $(50)$ are used during evaluation. Using $\lambda>0$ gives best results when fewer diffusion steps are used $(2,16)$. We note that the performance slightly deteriorates with $\lambda>0$ when the number of diffusion steps is large, see Figure 9. See Figure 19 for similar results on other Libero suites.
@@ Appendix A Notation
Name Notation Definition Appeared 1st
$\cdot$ $\rho$ Continuous negative definite kernel Sec 2.2
$\cdot$ $k$ Continuous positive definite kernel Sec 2.2
Scoring rule $S(p,y)$cdot Sec 2.2
Expected score $S(p,q)$ $S(p,q)=\mathbb{E}_{q}[S(p,Y)]$ Sec 2.2
Kernel score $S_{\rho}(p,y)$ $S_{\rho}(p,y)=\frac{1}{2}\mathbb{E}[\rho(X,X^{\prime})]-\mathbb{E}[\rho(X,y)]$ Sec 2.2
Energy score $S_{\beta}(p,y)$ $S_{\beta}(p,y)=S_{\rho}(p,y)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Sec 2.2
Expected Kernel score $S_{\rho}(p,q)$ $S_{\rho}(p,q)=\mathbb{E}_{q}[S_{\rho}(p,Y)]$ Sec 2.2
Expected Energy score $S_{\beta}(p,q)$ $S_{\beta}(p,q)=S_{\rho}(p,q)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Sec 2.2
Generalized (Gen.) Kernel score $S_{\lambda,\rho}(p,y)$ $S_{\lambda,\rho}(p,y)=\frac{\lambda}{2}\mathbb{E}_{p\otimes p}[\rho(X,X^{\prime})]-\mathbb{E}[\rho(X,y)]$ Sec 2.2
Generalized Energy score $S_{\lambda,\beta}(p,y)$ $S_{\lambda,\beta}(p,y)=S_{\lambda,\rho}(p,y)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Sec 2.2
Expected Gen. Kernel score $S_{\lambda,\rho}(p,q)$ $S_{\lambda,\rho}(p,q)=\mathbb{E}_{q}[S_{\lambda,\rho}(p,Y)]$ Sec 2.2
Expected Gen. Energy score $S_{\lambda,\beta}(p,q)$ $S_{\lambda,\beta}(p,q)=S_{\lambda,\rho}(p,q)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Sec 2.2
Squared MMD $D_{\rho}(p,q)$ $D_{\rho}(p,q)=S_{\rho}(q,q)-S_{\rho}(p,q)$ Sec 2.2
Energy Distance $D_{\beta}(p,q)$ $D_{\beta}(p,q)=D_{\rho}(p,q)$ with $\rho(x,x^{\prime})=\|x-x^{\prime}\|^{\beta}$ Sec 2.2
Diffusion Loss $\mathcal{L}_{\text{diff}}(\theta)$ $\mathcal{L}_{\text{diff}}(\theta)=\int^{1}_{0}w_{t}\mathbb{E}_{p_{0,t}}\left[||X_{0}-\hat{x}_{\theta}(t,X_{t})||^{2}\right]\mathrm{d}t$ Sec 2.1
Energy Diffusion Loss $\mathcal{L}(\theta)$ $\mathcal{L}(\theta)=-\int_{0}^{1}w_{t}\mathbb{E}_{p_{0,t}}\left[S_{\lambda,\beta}(p_{0|t}^{\theta}(\cdot|X_{t}),X_{0})\right]\mathrm{d}t$ Sec 3
Kernel Diffusion Loss $\mathcal{L}(\theta)$ $\mathcal{L}(\theta)=-\int_{0}^{1}w_{t}\mathbb{E}_{p_{0,t}}\left[S_{\lambda,\rho}(p_{0|t}^{\theta}(\cdot|X_{t}),X_{0})\right]\mathrm{d}t$ Sec 3
Joint Diffusion Energy Loss $\mathcal{L}_{\mathrm{joint}}(\theta)$ $\mathcal{L}_{\mathrm{joint}}(\theta)=-\int_{0}^{1}w_{t}\mathbb{E}_{p_{0,t}}\left[S_{\lambda,\beta}(p_{0|t}^{\theta}p_{t},(X_{0},X_{t}))\right]\mathrm{d}t$ Sec 4.2
@@ Appendix B Correspondence between Discrepancy and Diffusion Loss
Proposition B.1: Assume that $D_{\rho}(p,q)=\|\mathbb{E}_{p}[X]-\mathbb{E}_{q}[X]\|^{2}$ and that $p=\updelta_{\hat{x}_{\theta}(t,X_{t})}$ and $q=p_{0|t}(\cdot|X_{t})$. Then, we have $\mathbb{E}_{p_{t}}[D_{\rho}(p,q)]=\mathbb{E}_{p_{0,t}}[\|X_{0}-\hat{x}_{\theta}(t,X_{t})\|^{2}]-\mathbb{E}_{p_{0,t}}[\|X_{0}-\mathbb{E}[X_{0}|X_{t}]\|^{2}],$ (25) where the second term on the r.h.s. is independent of $\theta$.
@@ D.1 Energy diffusion loss and Maximum Mean Discrepancy diffusion loss
In this section, we first compare the energy diffusion loss (14) and another loss based on Maximum Mean Discrepancy (MMD), denoted MMD diffusion loss. We show that these two losses are equal up to a constant which does not depend on the network parameters $\theta$. However, this additional term appearing in the MMD diffusion loss is intractable, even though it could be potentially estimated using important sampling techniques. We consider the case where the kernel $k$ is given by $k(x,x^{\prime})=-\|x-x^{\prime}\|$. While our discussion can be extended to other settings, we also focus on the case where $\lambda=1,\beta=1$ and $w_{t}=1$ for simplicity.
@@ D.2 Joint Diffusion Energy Loss
We recall that the term which does not depend on $\theta$ in $\mathcal{L}_{\mathrm{MMD}}$ is difficult to approximate, the proposed approximation using importance sampling (74) is indeed expected to perform poorly. On the contrary, the term which does not depend on $\theta$ in $\mathcal{L}_{\mathrm{MMD},\mathrm{joint}}$ is very easy to approximate. This suggests that if one would optimize the loss with respect to the parameters of the kernel, then the choice $\mathcal{L}_{\mathrm{MMD},\mathrm{joint}}$ is more suitable. However, in the case where the parameters are fixed, as it is the case in our framework, then we will show in Section D.5 that $\mathcal{L}_{\mathrm{MMD}}$ is more sample efficient.
@@ D.5 Sample efficiency
Appendix E Diffusion-compatible kernels
@@ Appendix F DDIM updates from a Stochastic Differential Equation perspective
We first show that different choices of forward process yields different interpolation densities $p_{s|0,t}$ in Section F.1. With a careful choice of the forward process and appropriate parameters, we show that we can recover DDIM updates (Song et al., 2021a) in Section F.2.
@@ F.1 General forward processes
In this section, we consider general forward processes. The main property of all those forward processes $(\mathbf{X}_{t})_{t\in[0,1]}$ is that they will satisfy
$\mathbf{X}_{t}=\alpha_{t}\mathbf{X}_{0}+\sigma_{t}\mathbf{Z}^{\prime},\qquad\mathbf{X}_{0}\sim p_{0},$ (140)
with $\mathbf{Z}^{\prime}\sim\mathcal{N}(0,\mathrm{Id})$. First, we define
@@
We also consider $\mathbf{Z}\sim\mathcal{N}(0,\mathrm{Id})$ and introduce the dynamics $(\mathbf{X}_{t})_{t\in[0,1]}$
$\mathrm{d}\mathbf{X}_{t}=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\frac{g_{t}^{2}}{2\sigma_{t}}(1-\varepsilon^{2})\mathbf{Z}\right]\mathrm{d}t+\varepsilon g_{t}\mathrm{d}\mathbf{B}_{t},$ (142)
where $(\mathbf{B}_{t})_{t\in[0,1]}$ is a $d$-dimensional Brownian motion. In (142), we identify different forward processes which are all non-Markov except in the case where $\varepsilon=1$. The remarkable property of these forward processes is that they all recover the interpolation (140). This means that the forward trajectories might be very different but they all admit the same marginal distributions. More formally, we get the following result.
Proposition F.1: For any $t\in[0,1]$, let $\varepsilon_{t}\in[0,1]$. Let $\mathbf{Z}\sim\mathcal{N}(0,\mathrm{Id})$ and $\mathbf{X}_{0}\sim p_{0}$. Additionally, assume that the following SDE admits a solution $\mathrm{d}\mathbf{X}_{t}=\left[\partial_{t}\log(\alpha_{t})\mathbf{X}_{t}+\frac{g_{t}^{2}}{2\sigma_{t}}(1-\varepsilon^{2})^{1/2}\mathbf{Z}\right]\mathrm{d}t+\varepsilon g_{t}\mathrm{d}\mathbf{B}_{t},$ (143) where $(\mathbf{B}_{t})_{t\in[0,1]}$ is a $d$-dimensional Brownian motion, $p(z)=\mathcal{N}(z;0,\mathrm{Id})$ and $p_{t}(\cdot|z)$ is the density of $\mathbf{X}_{t}$ conditionally to $\mathbf{Z}=z$. Denote $p_{t}(x_{t})=\int_{\mathbb{R}^{d}}p_{t}(x_{t}|z)p(z)\mathrm{d}z$. We have that for any $t\in[0,1]$, $\alpha_{t}\mathbf{X}_{0}+\sigma_{t}\mathbf{Z}\sim p_{t}$.
@@
$\mathbf{X}_{t}=\frac{\alpha_{t}}{\alpha_{s}}\mathbf{X}_{s}+(1-\varepsilon^{2})^{1/2}\left(\sigma_{t}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}\right)\mathbf{Z}+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}^{2}}{\alpha_{s}^{2}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}},$ (144)
@@
$\mathbf{X}_{t}-\alpha_{t}\mathbf{X}_{0}$ $=\frac{\alpha_{t}}{\alpha_{s}}(\mathbf{X}_{s}-\alpha_{s}\mathbf{X}_{0})+(1-\varepsilon^{2})^{1/2}\left(\sigma_{t}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}\right)\mathbf{Z}+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}^{2}}{\alpha_{s}^{2}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}}$ (145)
$=\frac{\alpha_{t}}{\alpha_{s}}[(1-\varepsilon^{2})^{1/2}\sigma_{s}\mathbf{Z}+\varepsilon\sigma_{s}\hat{\mathbf{Z}}]+(1-\varepsilon^{2})^{1/2}\left(\sigma_{t}-\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}\right)\mathbf{Z}+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}^{2}}{\alpha_{s}^{2}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}}$ (146)
$=(1-\varepsilon^{2})^{1/2}\sigma_{t}\mathbf{Z}+\varepsilon\sigma_{s}\frac{\alpha_{t}}{\alpha_{s}}\hat{\mathbf{Z}}+\varepsilon\left(\sigma_{t}^{2}-\frac{\alpha_{t}^{2}}{\alpha_{s}^{2}}\sigma_{s}^{2}\right)^{1/2}\tilde{\mathbf{Z}},$ (147)
with $\mathbf{X}_{s}=\alpha_{s}\mathbf{X}_{0}+\sigma_{s}(1-\varepsilon)^{1/2}\mathbf{Z}+\sigma_{s}\varepsilon\hat{\mathbf{Z}}$, with $\mathbf{Z},\tilde{\mathbf{Z}}$ and $\hat{\mathbf{Z}}$ independent Gaussian random variables with zero mean and identity covariance. We can now compute the probability distribution $p_{s,t|0}$. This will allow us to compute $p_{s|0,t}$ using Gaussian posteriors. Using (147), we have